# Weak lensing of bright standard sirens: prospects for $\sigma_{8}$

### v1

@@ Abstract
I Introduction
@@
Gravitational waves (GWs) from compact binary coalescences provide a unique avenue to probe cosmic expansion through the construction of a Hubble diagram that is independent of the traditional cosmic distance ladder. GW events with electromagnetic counterparts act as bright standard sirens, allowing direct measurements of luminosity distances from the waveform amplitude. While standard candles have established the accelerated expansion of the Universe [1, 2], bright sirens offer independent constraints on the Hubble constant and other cosmological parameters [3, 4, 5, 6].
@@
The propagation of GWs across cosmological distances is affected by intervening cosmic structures through gravitational lensing. Searches for such signatures are actively underway within the LIGO-Virgo-KAGRA (LVK) network. Analyses from the fourth observing run have reported no definitive detections, although an intriguing candidate, GW231123, has been identified [7, 8, 9]. These early studies highlight the increasing importance of accounting for lensing systematics as detector sensitivities improve and event catalogues grow.
@@
In the weak-lensing regime, gravitational lensing results in the magnification or demagnification of the observed strain by a factor of $\sqrt{\mu}$ relative to the unlensed signal. The resulting scatter in luminosity distances is characterized by the magnification probability distribution ${\rm d}P/{\rm d}\mu$, which is non-Gaussian and evolves with redshift. At the redshifts probed by third-generation ground-based detectors such as ET [10] and space-based detectors such as LISA [11], weak-lensing scatter will contribute percent-level uncertainties to individual distance measurements and must therefore be incorporated carefully into cosmological analyses.
@@
Accurate modeling of ${\rm d}P/{\rm d}\mu$ is essential for robust cosmological inference with bright sirens, as simplified treatments can bias parameter estimates and underestimate true uncertainties [12, 13, 14, 15, 16]. In addition, ${\rm d}P/{\rm d}\mu$ encodes information about matter inhomogeneities. Measurements of the scatter in the Hubble diagram can thus provide insights into cosmology and the properties of cosmic structures. In particular, the scatter is sensitive to the standard deviation of matter perturbations, $\sigma_{8}$. The potential of lensing-induced scatter to constrain $\sigma_{8}$ has been explored for standard candles in [17, 18]. Recent results from the Dark Energy Survey demonstrate a $5\sigma$ detection of weak-lensing magnification of Type Ia supernovae [19], confirming that Hubble diagram scatter can indeed serve as a probe of $\sigma_{8}$.
@@
The possibility of constraining $\sigma_{8}$ through weak lensing of bright sirens was first explored in [20], where the lensing signal was characterized via the convergence power spectrum computed at the linear level. This approach was subsequently extended to a larger set of cosmological parameters and dark sirens in [21], and to cross-correlations with galaxy fields in [22]. The linear convergence power spectrum is valid only on large scales, where a few hundred sources are required for a reliable measurement. With a sample of $\mathcal{O}(10^{4})$ sources, a $3\%$ measurement of $\sigma_{8}$ was predicted in [20].
@@
In this work, we instead exploit the full magnification distribution, incorporating modeling of non-linear structures. This approach, first mentioned in [20], is complementary to that based on the convergence power spectrum and captures additional information, including lensing of individual sources and non-Gaussianity, allowing better sensitivity to $\sigma_{8}$. Indeed, we find that ET could measure $\sigma_{8}$ to $10\%$ with $300$ sources, and LISA could reach $30\%$ with just $12$ sources. Modeling non-linear structures, however, introduces theoretical uncertainties [16, 23], and further work beyond the scope of this study is needed to refine the magnification distribution. The goal of this work is to demonstrate the prospects of this approach by implementing, for the first time, the magnification distribution in Hubble-diagram reconstruction. In addition to $\sigma_{8}$, we consider the Hubble parameter $h$ and the matter abundance $\Omega_{M}$, that primarily affect the mean luminosity distance-redshift relation.
II Lensing magnification
@@
We compute the lensing magnification distribution ${\rm d}P/{\rm d}\mu$ using the stochastic approach developed in [24, 25, 26], which we have implemented in C++. [^1] The full code to generate the magnification distribution and reconstruct the Hubble diagram is publicly available at https://github.com/vianvask/halos. In this approach, the matter distribution along the line of sight is modeled as an ensemble of discrete lenses drawn from a specified mass function. For each realization, lenses are placed in redshift slices near the line of sight to the source, and the total magnification is obtained by summing the individual lensing contributions under the weak-lensing approximation. Repeating this procedure over many realizations yields the magnification probability distribution. This method is computationally efficient and significantly faster than approaches based on ray tracing through universes generated, for example, with N-body simulations [27, 28, 29, 30].
@@
The lensing magnification is given by the convergence $\kappa$ and the shear $\gamma=\sqrt{\gamma_{1}^{2}+\gamma_{2}^{2}}$ as
@@
In the weak-lensing approximation, $\kappa$, $\gamma_{1}$ and $\gamma_{2}$ of multiple lenses are obtained by summing the contributions of individual lenses: $\kappa=\sum_{j}\kappa_{\epsilon}(\boldsymbol{\theta}_{j})$, $\gamma_{1}=\sum_{j}\gamma_{\epsilon}(\boldsymbol{\theta}_{j})\cos\phi_{j}$ and $\gamma_{2}=\sum_{j}\gamma_{\epsilon}(\boldsymbol{\theta}_{j})\sin\phi_{j}$, where $\phi_{j}$ denotes the polar angle of the lens in the lens plane, defined with respect to a coordinate system common to all lenses, and $\boldsymbol{\theta}_{j}$ encodes the lens redshift $z_{j}$, its projected distance $r_{j}$ from line of sight to the source, and parameters describing the lens properties and orientation. We include lensing contributions from both halos and filaments.
We model halos using the pseudo-elliptical Navarro-Frenk-White (NFW) profile, for which the expressions for $\kappa$ and $\gamma$ are derived in [31]. The ellipticity is defined as $\epsilon=1-b/a$, where $a$ and $b$ are the semi-major and semi-minor axes of the lensing potential projected onto the lens plane. The ellipticity $\epsilon$, the scale radius $r_{s}$, and the characteristic density $\rho_{s}$ at the scale radius are functions of the halo mass $M$. The halo mass is defined as $M=(4\pi/3)\,\Delta\,\rho_{c}\,r_{\Delta}^{3}$, where $r_{\Delta}=C\,r_{s}$ and $\rho_{c}$ denotes the critical density. We adopt $\Delta=200$, use the concentration-mass relation $C=C(M,z)$ from [32], and the ellipticity fit $\epsilon=\epsilon(M,z)$ from [33]. Consequently, halos are characterized by the parameter set $\boldsymbol{\theta}_{j}=\{z_{j},r_{j},M_{j},\phi_{e,j}\}$, where $\phi_{e,j}$ denotes the orientation of the halo ellipticity.
We model filaments as cylindrical structures with a radial density profile $\rho\propto(1+r^{2}/r_{s}^{2})^{-1}$, where $r_{s}$ denotes the filament scale radius. The filament mass is defined as $M=\pi r_{s}^{2}L\,\bar{\rho}(r<r_{s})/2$, where $L$ is the filament length and $\bar{\rho}(r<r_{s})$ is the average density within radius $r_{s}$. Motivated by the results of [34], we fix $r_{s}=1\,\mathrm{Mpc}\,(M/10^{14}\,M_{\odot})^{1/3}$, $L=20\,\mathrm{Mpc}\,(M/10^{14}\,M_{\odot})^{1/3}$, and $\bar{\rho}(r<r_{s})=10\,\rho_{c}$. Filaments are characterized by the set of parameters $\boldsymbol{\theta}_{j}=\{z_{j},r_{j},M_{j},\theta_{f,j},\phi_{f,j}\}$, where the angles $\theta_{f,j}$ and $\phi_{f,j}$ determine the orientation of the filament with respect to the lens plane.
@@
We estimate the halo and filament mass functions using the excursion-set formalism [35], adopting the mass-dependent collapse barriers presented in [36]. The average comoving number density of potential lenses is given by
@@
where $\rho_{m}$ denotes the average present-day mass density and $A=[1+2^{-p}\Gamma(1/2-p)/\sqrt{\pi}]^{-1}$ ensures normalization. We adopt $\{p,q\}=\{0.3,0.8\}$ for halos and $\{p,q\}=\{0,0.7\}$ for filaments. The parameter $\nu$ is defined as $\nu\equiv\delta_{c}(z)^{2}/\sigma_{M}^{2}$, where $\delta_{c}(z)$ is the density contrast threshold for spherical collapse and $\sigma_{M}^{2}$ is the variance of the matter fluctuations at the mass scale $M$. We compute the latter using the real space top-hat window function and the matter power spectrum estimated as in [37]. We fix the normalization by setting the value of $\sigma_{8}=\sigma_{M}(R=8h\,{\rm Mpc})$.
We also include the linear bias factor $B(M_{b},z)$ (see, e.g. [38]). The halo number density is modeled as a biased lognormal field ${\rm d}n={\rm d}\bar{n}\exp[B(M_{b},z)\delta_{b}-B(M_{b},z)^{2}\sigma_{M_{b}}^{2}/2]$ where $\delta_{b}$ is a Gaussian random variable with variance $\sigma_{M_{b}}^{2}$ and the mass scale $M_{b}$ corresponds to the average mass enclosed within a radius $r_{\rm max}$ defined below and within a redshift bin $\Delta z$, chosen so that $M_{b}$ is much larger than the typical lens masses.
@@
with $H_{0}=100h\,{\rm km}/{\rm s}/{\rm Mpc}$. We impose an upper limit $r_{\rm max}=r_{\rm max}(M,z)$ on the transverse distance $r$ of the lens from the line of sight to the source by requiring that for each individual lens $\kappa$ exceeds a threshold value $\kappa_{\rm thr}$ that we choose so that the average number of lens halos is $\bar{N}_{l,h}(r<r_{\rm max})=100$. In Fig. 1 we show the PDF of the lenses for the sources at $z_{s}=5$, integrated over $r$. The threshold in this case is $\kappa_{\rm thr}\approx 0.0007$ and lenses in the purple region would give $\kappa<\kappa_{\rm thr}$.
We have not taken into account the wave optics effects that suppress the magnification for [39, 40]
@@
We compute the distribution of lensing magnifications, $\mu$, by generating realizations of halos and filaments along the line of sight to a source at redshift $z_{s}$. For each realization $i$ and redshift bin, we draw the number of lenses from a Poisson distribution whose mean we compute accounting for the linear bias. We then sample the lens parameters from the probability distribution function in Eq. (3) and calculate the corresponding convergence $\kappa_{i}$ and shear $\gamma_{i}$. Next, we subtract the ensemble mean of $\{\kappa_{i}\}$ from each realization and compute the magnification $\mu_{i}$. We bin the magnifications and apply a $1/\mu$ factor to convert the image-plane distribution into the source-plane distribution. The resulting probability distribution functions of the lensing magnification for different source redshifts $z_{s}$, obtained from $4\times 10^{5}$ realizations at each $z_{s}$, are shown in Fig. 2 for the cosmological parameters inferred from CMB observations ($h=0.674$, $\Omega_{M}=0.315$, $\sigma_{8}=0.811$[41], which we adopt as our benchmark cosmology.
@@
In Fig. 3, we show the variance of apparent luminosity distances as a function of redshift. The open symbols show the results obtained by sampling $\mu$ from the magnification PDFs presented in Fig. 2. The curves show the fitting function [42]
@@
Our full model, including elliptical halos, bias and filaments, gives $c=0.22$, $\beta=1.09$, $\alpha=1.87$. We find reasonably good agreement with the results of [29], which were obtained via ray-tracing through universes generated with high-resolution $N$-body simulations and gives $c=0.16$ $\beta=0.73$, $\alpha=1.91$.
Figure 3: The variance of luminosity distances induced by weak lensing in various approximations of the lens population. For comparison, the dashed curve shows the result of [29]. The open symbols depict the numerical results, while the curves illustrate the fit using the ansatz (8).
@@
III Hubble diagram reconstruction
We generate benchmark catalogues similar to those considered in previous studies (e.g. [43, 16]). Specifically, we construct one catalogue consisting of $300$ binary neutron star (BNS) systems and another consisting of $12$ binary massive black hole (BMBH) systems. The luminosity distances of the BNS population are assumed to be measurable from the GW signals by ET, while those of the BMBH population are assumed to be measurable by LISA. The redshifts are assumed to be obtained from electromagnetic counterparts. BNS mergers are expected to produce multiple types of detectable electromagnetic counterparts, as demonstrated by GW170817 [5]. In contrast, electromagnetic counterparts of BMBHs remain uncertain. However, several studies suggest that a small number of such events may be detectable [44, 45, 46, 43].

### v2

@@ Abstract
keywords:
gravitational waves – weak lensing – cosmological parameters
pubyear: 2026volume: 547pagerange: Weak lensing of bright standard sirens: prospects for $\sigma_{8}$References
1 Introduction
@@
Gravitational waves (GWs) from compact binary coalescences provide a unique avenue to probe cosmic expansion through the construction of a Hubble diagram that is independent of the traditional cosmic distance ladder. GW events with electromagnetic (EM) counterparts act as bright standard sirens, allowing direct measurements of luminosity distances from the waveform amplitude. While standard candles have established the accelerated expansion of the Universe (Riess et al., 1998; Perlmutter et al., 1999), bright sirens offer independent constraints on the Hubble constant and other cosmological parameters (Schutz, 1986; Holz & Hughes, 2005; Abbott et al., 2017; Jin et al., 2026).
@@
The propagation of GWs across cosmological distances is affected by intervening cosmic structures through gravitational lensing. Searches for such signatures are actively underway within the LIGO-Virgo-KAGRA (LVK) network. Analyses from the fourth observing run have reported no definitive detections, although an intriguing candidate, GW231123, has been identified (Chan et al., 2025; Goyal et al., 2025; Chakraborty & Mukherjee, 2025; Shan et al., 2025). These early studies highlight the increasing importance of accounting for lensing systematics as detector sensitivities improve and event catalogues grow.
@@
In the weak-lensing regime, gravitational lensing results in the magnification or demagnification of the observed strain by a factor of $\sqrt{\mu}$ relative to the unlensed signal. The resulting scatter in luminosity distances is characterized by the magnification probability distribution ${\rm d}P/{\rm d}\mu$, which is non-Gaussian and evolves with redshift. At the redshifts probed by third-generation ground-based detectors such as ET (Abac et al., 2025a) and space-based detectors such as LISA (Amaro-Seoane et al., 2017), weak-lensing scatter will contribute percent-level uncertainties to individual distance measurements and must therefore be incorporated carefully into cosmological analyses.
@@
Accurate modeling of ${\rm d}P/{\rm d}\mu$ is essential for robust cosmological inference with bright sirens, as simplified treatments can bias parameter estimates and underestimate true uncertainties (Fleury et al., 2017; Adamek et al., 2019; Cusin & Tamanini, 2021; Canevarolo & Chisari, 2024; Mpetha et al., 2024). In addition, ${\rm d}P/{\rm d}\mu$ encodes information about matter inhomogeneities. Measurements of the scatter in the Hubble diagram can thus provide insights into cosmology and the properties of cosmic structures. In particular, the scatter is sensitive to the standard deviation of matter perturbations, $\sigma_{8}$. The potential of lensing-induced scatter to constrain $\sigma_{8}$ has been explored for standard candles by Quartin et al. (2014); Castro & Quartin (2014). Recent results from the Dark Energy Survey demonstrate a $5\sigma$ detection of weak-lensing magnification of Type Ia supernovae (Shah et al., 2024), confirming that Hubble diagram scatter can indeed serve as a probe of $\sigma_{8}$.
@@
The possibility of constraining $\sigma_{8}$ through weak lensing of bright sirens was first explored by Congedo & Taylor (2019), where the lensing signal was characterized via the convergence power spectrum computed at the linear level. This approach was subsequently extended to a larger set of cosmological parameters and dark sirens by Mpetha et al. (2023), and to cross-correlations with galaxy fields by Balaudo et al. (2023). The linear convergence power spectrum is valid only on large scales, where a few hundred sources are required for a reliable measurement. With a sample of $\mathcal{O}(10^{4})$ sources, a $3\%$ measurement of $\sigma_{8}$ was predicted by Congedo & Taylor (2019).
@@
In this work, we instead exploit the full magnification distribution, incorporating modeling of non-linear structures. This approach, first mentioned by Congedo & Taylor (2019), is complementary to that based on the convergence power spectrum and captures additional information, including lensing of individual sources and non-Gaussianity, allowing better sensitivity to $\sigma_{8}$. Indeed, we find that ET could measure $\sigma_{8}$ to $10\%$ with $300$ sources, and LISA could reach $30\%$ with just $12$ sources. Modeling non-linear structures, however, introduces theoretical uncertainties (Mpetha et al., 2024; Alfradique et al., 2024), and further work beyond the scope of this study is needed to refine the magnification distribution. The goal of this work is to demonstrate the prospects of this approach by implementing, for the first time, the magnification distribution in Hubble-diagram reconstruction. In addition to $\sigma_{8}$, we consider the Hubble parameter $h$ and the matter abundance $\Omega_{M}$, that primarily affect the mean luminosity distance-redshift relation.
2 Lensing magnification
@@
We compute the lensing magnification distribution ${\rm d}P/{\rm d}\mu$ using the stochastic approach developed by Kainulainen & Marra (2009, 2011a, 2011b), which we have implemented in C++.111 The full code to generate the magnification distribution and reconstruct the Hubble diagram is publicly available at https://github.com/vianvask/halos. In this approach, the matter distribution along the line of sight is modeled as an ensemble of discrete lenses drawn from a specified mass function. For each realization, lenses are placed in redshift slices near the line of sight to the source, and the total magnification is obtained by summing the individual lensing contributions under the weak-lensing approximation. Repeating this procedure over many realizations yields the magnification probability distribution. This method is computationally efficient and significantly faster than approaches based on ray tracing through universes generated, for example, with N-body simulations (Holz & Wald, 1998; Holz & Linder, 2005; Takahashi et al., 2011; Takahashi et al., 2017). Furthermore, it has been argued e.g. by Shah et al. (2025) that N-body simulations do not have sufficient resolution to accurately reconstruct the lensing magnification distribution for point sources.
@@
The lensing magnification, defined as the ratio of the lensed to unlensed flux, is given by the convergence $\kappa$ and the shear $\gamma=\sqrt{\gamma_{1}^{2}+\gamma_{2}^{2}}$ as
@@
In the weak-lensing approximation, $\kappa$, $\gamma_{1}$ and $\gamma_{2}$ of multiple lenses are obtained by summing the contributions of individual lenses:222 We retain the non-linear relation (1) to capture high-magnification events, which are dominated by single lenses. Our approach smoothly connects this regime to low magnifications, where linear superposition of $\kappa$ and $\gamma_{1,2}$ is adequate. $\kappa=\sum_{j}\kappa_{\epsilon}(\boldsymbol{\theta}_{j})$, $\gamma_{1}=\sum_{j}\gamma_{\epsilon}(\boldsymbol{\theta}_{j})\cos\phi_{j}$ and $\gamma_{2}=\sum_{j}\gamma_{\epsilon}(\boldsymbol{\theta}_{j})\sin\phi_{j}$, where $\phi_{j}$ denotes the polar angle of the lens in the lens plane, defined with respect to a coordinate system common to all lenses, and $\boldsymbol{\theta}_{j}$ encodes the lens redshift $z_{j}$, its projected distance $r_{j}$ from line of sight to the source, and parameters describing the lens properties and orientation. We include lensing contributions from both halos and filaments.
We model halos using the pseudo-elliptical Navarro-Frenk-White (NFW) profile, for which the expressions for $\kappa$ and $\gamma$ are derived by Golse & Kneib (2002). The ellipticity is defined as $\epsilon=1-b/a$, where $a$ and $b$ are the semi-major and semi-minor axes of the lensing potential projected onto the lens plane. The ellipticity $\epsilon$, the scale radius $r_{s}$, and the characteristic density $\rho_{s}$ at the scale radius are functions of the halo mass $M$. The halo mass is defined as $M=(4\pi/3)\,\Delta\,\rho_{c}\,r_{\Delta}^{3}$, where $r_{\Delta}=C\,r_{s}$ and $\rho_{c}$ denotes the critical density. We adopt $\Delta=200$, use the concentration-mass relation $C=C(M,z)$ from Dutton & Macciò (2014), and the ellipticity fit $\epsilon=\epsilon(M,z)$ from Allgood et al. (2006). Consequently, halos are characterized by the parameter set $\boldsymbol{\theta}_{j}=\{z_{j},r_{j},M_{j},\phi_{e,j}\}$, where $\phi_{e,j}$ denotes the orientation of the halo ellipticity.
We model filaments as cylindrical structures with a radial density profile $\rho\propto(1+r^{2}/r_{s}^{2})^{-1}$, where $r_{s}$ denotes the filament scale radius. The filament mass is defined as $M=\pi r_{s}^{2}L\,\bar{\rho}(r<r_{s})/2$, where $L$ is the filament length and $\bar{\rho}(r<r_{s})$ is the average density within radius $r_{s}$. Motivated by the results of Colberg et al. (2005), we fix $r_{s}=1\,\mathrm{Mpc}\,(M/10^{14}\,M_{\odot})^{1/3}$, $L=20\,\mathrm{Mpc}\,(M/10^{14}\,M_{\odot})^{1/3}$, and $\bar{\rho}(r<r_{s})=10\,\rho_{c}$. Filaments are characterized by the set of parameters $\boldsymbol{\theta}_{j}=\{z_{j},r_{j},M_{j},\theta_{f,j},\phi_{f,j}\}$, where the angles $\theta_{f,j}$ and $\phi_{f,j}$ determine the orientation of the filament with respect to the lens plane.
@@
We estimate the halo and filament mass functions using the excursion-set formalism (Bond et al., 1991), adopting the mass-dependent collapse barriers presented by Shen et al. (2006). The average comoving number density of potential lenses is given by
@@
where $\rho_{m}$ denotes the average present-day mass density and $A=[1+2^{-p}\Gamma(1/2-p)/\sqrt{\pi}]^{-1}$ ensures normalization. We adopt $\{p,q\}=\{0.3,0.8\}$ for halos and $\{p,q\}=\{0,0.7\}$ for filaments.333 The values of $p$ and $q$ were determined by performing random walks and fitting the first-crossing distribution, following Sheth & Tormen (2002). The parameter $\nu$ is defined as $\nu\equiv\delta_{c}(z)^{2}/\sigma_{M}^{2}$, where $\delta_{c}(z)$ is the density contrast threshold for spherical collapse and $\sigma_{M}^{2}$ is the variance of the matter fluctuations at the mass scale $M$. We compute the latter using the real space top-hat window function and the matter power spectrum estimated as Eisenstein & Hu (1998). We fix the normalization by setting the value of $\sigma_{8}=\sigma_{M}(R=8h\,{\rm Mpc})$.
We also include the linear bias factor $B(M_{b},z)=1+(q\nu-1)/\delta_{c}(0)+2p/[\delta_{c}(0)(1+(q\nu)^{p})]$ with $\{p,q\}=\{0.3,0.75\}$ (see, e.g. Baumann (2022)). The halo number density is modeled as a biased lognormal field ${\rm d}n={\rm d}\bar{n}\exp[B(M_{b},z)\delta_{b}-B(M_{b},z)^{2}\sigma_{M_{b}}^{2}/2]$ where $\delta_{b}$ is a Gaussian random variable with variance $\sigma_{M_{b}}^{2}$ and the mass scale $M_{b}$ corresponds to the average mass enclosed within a radius $r_{\rm max}$ defined below and within a redshift bin $\Delta z$, chosen so that $M_{b}$ is much larger than the typical lens masses.
@@
with $H_{0}=100h\,{\rm km}/{\rm s}/{\rm Mpc}$. We impose an upper limit $r_{\rm max}=r_{\rm max}(M,z)$ on the transverse distance $r$ of the lens from the line of sight to the source by requiring that for each individual lens $\kappa$ exceeds a threshold value $\kappa_{\rm thr}$ that we choose so that the average number of lens halos is $\bar{N}_{l,h}(r<r_{\rm max})=100$.444 This choice is made to balance computational efficiency and accuracy. In Fig. 1 we show the PDF of the lenses for the sources at $z_{s}=5$, integrated over $r$. The threshold in this case is $\kappa_{\rm thr}\approx 0.0007$ and lenses in the purple region would give $\kappa<\kappa_{\rm thr}$.
We have not taken into account the wave optics effects that suppress the magnification for (Fairbairn et al., 2023; Urrutia & Vaskonen, 2025)
@@
We compute the distribution of lensing magnifications, $\mu$, by generating realizations of halos and filaments along the line of sight to a source at redshift $z_{s}$. For each realization $i$ and redshift bin, we draw the number of lenses from a Poisson distribution whose mean we compute accounting for the linear bias. We then sample the lens parameters from the probability distribution function in Eq. (3) and calculate the corresponding convergence $\kappa_{i}$ and shear $\gamma_{i}$. Next, we subtract the ensemble mean of $\{\kappa_{i}\}$ from each realization and compute the magnification $\mu_{i}$. We bin the magnifications and apply a $1/\mu$ factor to convert the image-plane distribution into the source-plane distribution. The resulting probability distribution functions of the lensing magnification for different source redshifts $z_{s}$, obtained from $4\times 10^{5}$ realizations at each $z_{s}$, are shown in Fig. 2 for the cosmological parameters inferred from CMB observations ($h=0.674$, $\Omega_{M}=0.315$, $\sigma_{8}=0.811$(Aghanim et al., 2020), which we adopt as our benchmark cosmology.
@@
In Fig. 3, we show the variance of apparent luminosity distances as a function of redshift.555 We note that the weak lensing variance is often given in terms of apparent magnitude $m=5\log_{10}D_{L}+{\rm const.}$ (see e.g. Marra et al. (2013)). The relation between weak lensing scatter in $m$ and in $D_{L}$ is given by $\sigma_{D_{L}}/D_{L}\approx 0.46\sigma_{m}$. The open symbols show the results obtained by sampling $\mu$ from the magnification PDFs presented in Fig. 2. The curves show the fitting function (Hirata et al., 2010)
@@
Our full model, including elliptical halos, bias and filaments, gives $c=0.22$, $\beta=1.09$, $\alpha=1.87$. We find reasonably good agreement with the results of Takahashi et al. (2011), which were obtained via ray-tracing through universes generated with high-resolution $N$-body simulations and gives $c=0.16$ $\beta=0.73$, $\alpha=1.91$.
Figure 3: The variance of luminosity distances induced by weak lensing in various approximations of the lens population. For comparison, the dashed curve shows the result of Takahashi et al. (2011). The open symbols depict the numerical results, while the curves illustrate the fit using the ansatz (8).
@@
3 Hubble diagram reconstruction
We generate benchmark catalogues similar to those considered in previous studies (e.g. Mangiagli et al. (2025); Mpetha et al. (2024)). Specifically, we construct one catalogue consisting of $300$ binary neutron star (BNS) systems and another consisting of $12$ binary massive black hole (BMBH) systems. The luminosity distances of the BNS population are assumed to be measurable from the GW signals by ET, while those of the BMBH population are assumed to be measurable by LISA. The redshifts are assumed to be obtained from EM counterparts. BNS mergers are expected to produce multiple types of detectable EM counterparts, as demonstrated by GW170817 (Abbott et al., 2017). Finding $\mathcal{O}(100)$ bright BNS events over a 10 year period is considered realistic (Iacovelli et al., 2022; Loffredo et al., 2025). In contrast, EM counterparts of BMBHs remain uncertain. However, several studies suggest that a small number of such events may be detectable (Dotti et al., 2006; Tamanini et al., 2016; Mangiagli et al., 2022; Mangiagli et al., 2025).